two spheres approach each other and collide elastically. If the masses are equal and the speeds are v1=3.0m/s and v2=-5.0m/s. Determine the final speeds of each
Answer #1

given

masses are equal

m1 = m2 = m

v1 = 3.0 m/sec

and

v2 = - 5.0 m/sec

the final speed is vf = ?

using the conservation of momentum

because the two spheres approach each other and collide elastically

m1 v1 + m2 v2 = ( m1 + m2 ) vf

m x 3 - m x 5 = ( m + m ) x vf

m ( 3 - 5 ) = 2 m x vf

- 2 = 2 x vf

vf = - 1 m/sec

so the final speeds of each sphere is vf = - 1 m/sec

here negative sign shows the direction.